### v1

# Incentive Compatibility for AI Alignment in Sociotechnical Systems: Positions and Prospects

Abstract
The burgeoning integration of artificial intelligence (AI) into human society brings forth significant implications for societal governance and safety. While considerable strides have been made in addressing AI alignment challenges, existing methodologies primarily focus on technical facets, often neglecting the intricate sociotechnical nature of AI systems, which can lead to a misalignment between the development and deployment contexts. To this end, we posit a new problem worth exploring: Incentive Compatibility Sociotechnical Alignment Problem (ICSAP). We hope this can call for more researchers to explore how to leverage the principles of Incentive Compatibility (IC) from game theory to bridge the gap between technical and societal components to maintain AI consensus with human societies in different contexts. We further discuss three classical game problems for achieving IC: mechanism design, contract theory, and Bayesian persuasion, in addressing the perspectives, potentials, and challenges of solving ICSAP, and provide preliminary implementation conceptions.
1 Introduction
“If we use, to achieve our purposes, a mechanical agency with whose operation we cannot interfere effectively $\dots$ we had better be quite sure that the purpose put into the machine is the purpose which we really desire.” (Wiener, 1960; Russell, 2019)
Norbert Wiener
The rapid development of artificial intelligence (AI) has had a significant impact on human society (Makridakis, 2017; Peeters et al., 2021; Wamba et al., 2021), from robots entering human production and living environments (Michaelis & Mutlu, 2018; Fu et al., 2024) to large language models (LLMs) capable of complex natural language interactions (Zhao et al., 2023; Bubeck et al., 2023) and reasoning ability (Wei et al., 2022; Wang et al., 2022). Therefore, an increasing number of people believe that as AI capabilities improve, AI systems will become integrated into human society in the future and be deployed in increasingly complex scenarios (Gladden, 2019; Dwivedi et al., 2021). Conversely, the powerful capabilities of AI systems have raised concerns about their safety (Cath et al., 2018; Peeters et al., 2021), especially considering their behavioral motivations  [^1] https://www.scai.gov.sg/scai-question-6/, alignment science  [^2] https://www.anthropic.com/news/core-views-on-ai-safety and how they align with human values and intentions (Ji et al., 2023). This is recognized as the “AI Alignment” problem.
Substantial progress has been made in addressing AI alignment issues, especially in the forward alignment process (Ji et al., 2023), which enables AI systems to have alignment capabilities 2. The methods for this process can mainly be divided into two categories. The first category involves learning from feedback (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), and there have been some significant research topics, including preference modeling (Akrour et al., 2011; Wirth et al., 2017), policy learning (Amodei et al., 2016; Ibarz et al., 2018), and scalable oversight (Christiano et al., 2018; Irving et al., 2018; Bai et al., 2022a; Burns et al., 2023). The second category focuses on resolving distributional shift (Di Langosco et al., 2022; Ngo et al., 2022) in learning, with notable subproblems including algorithmic interventions (Vapnik, 1991; Krueger et al., 2021; Lubana et al., 2023), adversarial training (Goodfellow et al., 2014; Poursaeed et al., 2021), and cooperative training (Dafoe et al., 2020, 2021).
However, these methods only consider the given alignment objectives, focusing solely on technical components such as dataset, architecture, and training algorithms, etc. (Weidinger et al., 2023b), overlooking the fact that AI systems are sociotechnical systems (Selbst et al., 2019). Some studies have indicated that relying solely on technical means will result in a sociotechnical gap between the model’s development context and its actual deployment context (Selbst et al., 2019; Lazar & Nelson, 2023), which is also detrimental to further social governance. Such examples are not uncommon in daily life. ChatGPT, trained on internet data and fine-tuned through RLHF (Ouyang et al., 2022), requires prompt engineering for adaptation to individual needs, highlighting unaddressed challenges in existing alignment techniques. Additionally, for sociotechnical systems, existing research is more concerned with only societal components like governance and evaluation methods (Raji et al., 2020; Dean et al., 2021; Weidinger et al., 2023b). Thus, currently, there are lack of means to simultaneously consider both technical and societal components, enabling AI systems themselves to maintain consensus with human society.
Incentive Compatibility (IC) (Hurwicz, 1972), derived from game theory, suggests that participants only need to pursue their true interests to reach optimal outcomes (Roughgarden, 2010). This concept leverages self-interested behavior, aligning actions with the game designer’s goals (Groves & Ledyard, 1987). With IC, each agent can maintain private goal information acquired during pretraining. Only by reconstructing different environments and rules,agents can optimize their own objectives to achieve outcomes that meet the needs of human society in different contexts. Therefore, we believe that exploring the IC property for AI alignment problems in sociotechnical systems is a highly worthwhile research endeavor.
In this paper, we separate a new subproblem from AI alignment problems in sociotechnical systems, called Incentive Compatibility Sociotechnical Alignment Problem (ICSAP), and based on this, we propose our main position:
Achieving incentive compatibility can simultaneously consider both technical and societal components in the forward alignment phase, enabling AI systems to keep consensus with human societies in different contexts.
2 Motivation and Opportunity: A Brief Example
In this section, we will use a very simple example to demonstrate how IC works in addressing AI alignment issues in sociotechnical systems through mechanism design which will be illustrated specifically in section 5.
Figure 1: A simple example illustrates how IC facilitates ICSAP scenarios through mechanism design. In the diagram, two agents aim to maximize cake consumption during technical training. However, the user desires equal cake distribution. Without IC, deploying both agents directly could lead to one party monopolizing the cake (a). With IC (b), the mechanism dictates that the second chooser is the one who cuts the cake. This ensures alignment with real-world needs by allowing agents to optimize within the rules, achieving the user’s goal and aligning sociotechnical systems.
Consider a classic divide and choose problem: the two-player cake cutting (Steinhaus, 1948). In this example (see Figure 1), two self-interested agents aim to maximize their cake share, while the human seeks an equal division. If either agent cuts the cake, they’ll take the whole. To align individual interests with the human’s goal, a simple mechanism is proposed: the cutter chooses second. This constraint ensures the agent’s pursuit of self-interest coincides with the host’s objective, achieving Alignment. The mechanism’s IC conditions facilitate consensus on an equal distribution and maximization of cake consumption.
If we consider the agent as an AI system and its desire to eat the most cake as the objective imparted by the technical component of training, we only need to use automated methods to search for corresponding mechanisms with IC properties as rules based on different real-world requirements to effectively solve ICSAP. Of course, hosts can also have different contextual needs, and they may not necessarily be self-interested. Here, we’re just using this one example to illustrate our point.
3 Related Work
To better understand the relevant content of ICSAP, in this section, we will introduce background work from two parts: AI alignment and sociotechnical systems.
AI Alignment.
The scope of AI alignment issues can mainly be divided into two processes, namely forward alignment and backward alignment (Ji et al., 2023).
The former refers to the development training and deployment training stages aimed at endowing AI systems with alignment capabilities. This process has yielded many significant achievements, including RLHF (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), DPO (Rafailov et al., 2023), etc., which align LLMs through fine-tuning methods; RLAIF (Bai et al., 2022b) RLHAIF (Cotra, 2021) IDA (Christiano et al., 2018) RRM (Leike et al., 2018; Hubinger, 2020) Debate (Irving et al., 2018; Michael et al., 2023) CIRL (Hadfield-Menell et al., 2016) Weak-To-Strong (Burns et al., 2023) and other scalable oversight methods; as well as REx (Krueger et al., 2021) CBFT (Lubana et al., 2023) and other works addressing distribution shift issues.
The latter mainly refers to the process of ensuring that AI systems can meet alignment-related requirements, including the assurance and governance phases. Current work includes ETHICS (Hendrycks et al., 2020), SOCIAL-CHEM-101 (Forbes et al., 2020), MIC (Ziems et al., 2022), TRUSTLLM (Sun et al., 2024), etc., which involve safety evaluation datasets. Methods for handling failure modes through adversarial attacks via red teaming (Perez et al., 2022; Liu et al., 2023); research on mechanistic interpretability (Olah et al., 2020) and governance criteria like ARC Evals  [^3] https://metr.org/blog/2023-03-18-update-on-recent-evals/ (Kinniment et al., 2023).
Sociotechnical Systems.
The concept of sociotechnical originates from labor studies in coal mining (Trist & Bamforth, 1951; Ropohl, 1999), where the technical components refer to optimizing components of machines themselves, such as data in AI systems, sampling methods, training algorithms, etc. (Weidinger et al., 2023b), while the societal components pertain to the complementary interaction between humans and machines as a social whole. Its core ideas are encapsulated in two principles  [^4] https://open.ncl.ac.uk/theories/9/socio-technical-theory/: Firstly, there is an inherent interaction and mutual influence between social and technical components in defining the success of a system. Secondly, there needs to be a certain degree of fit between social and technical aspects to achieve their joint optimization. Fundamentally, this implies equal consideration of technological and human factors throughout the entire process of societal technological design or redesign (Emery, 1980; Trist, 1981; Emery, 1993).
Lazar & Nelson (2023); Shelby et al. (2023) have revealed that considering AI systems solely from a technical component perspective is insufficient to ensure their safety. It is also necessary to consider the context of human society to ensure genuine safety in the practical deployment of AI systems. Existing research on alignment in sociotechnical systems primarily focuses on the evaluation of societal components. Selbst et al. (2019) identifies five challenges for AI technologies in sociotechnical systems, proposing a shift in design focus towards processes and considering social participants. Weidinger et al. (2023b) advocates for a three-tier evaluation framework to assess AI risks, emphasizing the need for both humans and machines in technology. They also propose methods to bridge safety assessment gaps. Dean et al. (2021) highlights the importance of considering dimensions such as optimization, consensus, value, and failure in AI research, exemplified by drones.
4 Background and Overview
In this section, we formulate the interactive dynamics among multiple AI and human populations coexisting in a societal context as a multi-player Markov game, referred to as Sociotechnical Game (SG). The game is assumed to take place in an infinite-horizon discounted setting. The complexity of this game can be extended by increasing the number of AI and humans, the frequency of interactions, and introducing elements such as incomplete information and imperfect information. Specific details will be adjusted based on the real-world scenarios and issues discussed later in the text. Here, we present a general formulation.
Formally, the SG is defined as a tuple $\mathcal{G}=(\mathcal{N}^{p},\mathcal{S}^{p},\mathcal{A}^{p},R^{p},\rho^{p},\gamma^{p},\mathbb{P})$ with the set of players $\mathcal{N}^{p}$, the state space $\mathcal{S}^{p}$, the joint action space $\mathcal{A}^{p}$, the reward function $R^{p}$, the initial state distribution $\rho^{p}$, the discount factor $\gamma^{p}$ and the transition probability function $\mathbb{P}$. We use $\mathcal{H}$ and $\mathcal{I}$ to represent human and AI, respectively. $p$ denotes a player. Given that the scenarios considered in this paper involve both human and AI, it is established that $p\in\{\mathcal{H},\mathcal{I}\}$. We use $n$ to denote the total number of players, where $n:=\left|\mathcal{H}^{p}\cup\mathcal{I}^{p}\right|$, In the more detailed definitions provided below, to ensure precision in the formulation, we will distinguish basic symbols in various contexts by incorporating superscripts or subscripts.
4.1 Basic Concepts in Human-AI Interaction
Policies. policy $\pi^{p}_{k}$ for the $k$-th player is a function mapping a given state to a distribution over available actions:
$\pi^{p}_{k}:\mathcal{S}^{p}\ni s\mapsto\pi^{p}_{k}(\cdot\mid s)\in\Delta\left(\mathcal{A}^{p}_{k}\right),$ (1)
where $k\in\mathbb{N}$ and $0<k\leq n$. Policies delineate how AI or human make decisions based on their current state, real-time observations, and the prevailing societal context. We let $\Pi^{p}:\mathcal{S}^{p}\rightarrow\Delta(\mathcal{A}^{p})$ denote the policy space for each player.
Value function. We use $V_{s}(\pi^{p}_{k},\pi^{p}_{-k})$ to denote the value function:
$V_{s}(\pi^{p}_{k},\pi^{p}_{-k}):\mathcal{S}^{p}\ni s,\Pi^{p}\ni{\pi}^{p}_{k}\mapsto\mathbb{R},$ (2)
where $-k$ denotes the set of player indices excluding the index corresponding to the $k$-th player. The value function is defined as the expected cumulative discounted reward at state $s\in\mathcal{S}^{p}$ under the joint policy $({\pi}^{p}_{k},{\pi}^{p}_{-k})$:
$V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{\left(\pi^{p}_{k},\pi^{p}_{-k}\right)}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{p}\left(s_{t},{a}^{k}_{t},a^{-k}_{t}\right)\mid s_{0}=s\right].\footnotesize$ (3)
The reward function $R^{p}$ encapsulates the short-term value of a particular strategy at a given moment. It provides a quantification of the immediate benefits or drawbacks associated with choosing a specific course of action. The initial state is sampled from the initial state distribution $\rho^{p}$. Therefore, the value function is denoted as:
$V_{\rho^{p}}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{s\sim\rho^{p}}\left[V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right)\right].$ (4)
The value function captures the enduring values of AI or humans, guiding policy selection in the current state. Estimating and controlling it is crucial for aligning AI with human values, enabling systematic evaluation and management of the long-term implications of current choices. This alignment fosters synergy between AI and human values.
4.2 Overview
Figure 2: The figure illustrates how IC tackles ICSAP based on three classic game-theoretic problems. Figure (a) illustrates the case of IC through mechanism design. The left side of the figure demonstrates a sociotechnical gap between agents considering only technical components and the values of real humans. On the right side, it shows that by designing corresponding mechanisms according to different needs, we can adjust the values of agents, aligning their utility with human requirements under IC conditions, thus achieving alignment in sociotechnical systems. For the left half of (b) and (c), both depict a sociotechnical gap between humans and a single agent. The right half of (b) demonstrates humans designing contracts that satisfy IC conditions based on specific needs, thereby adjusting the values of the agent through the contract. The right half of (c) illustrates a scenario of Bayesian persuasion where humans design information satisfying IC conditions according to their own needs, allowing agents to choose actions maximizing human demands without compromising their own values, thus solving ICSAP.
We consider that AI system agents, which take into account both human factors and only technical components during the development process, collectively form a sociotechnical scenario. We assume that these AI agents exhibit self-interested characteristics and will seek to maximize the value or objectives generated during their training process. In this scenario, we aim to demonstrate that IC can simultaneously address the technical components and societal components of AI systems, enabling AI to achieve consensus with human society in different contexts by adjusting different incentive conditions.
In the following sections, we will demonstrate three classic game problems as applying media of IC properties: Mechanism Design in section 5, Contract Theory in section 6, and Bayesian Persuasion in section 7. An overview of these approaches is depicted in Figure 2.
5 Mechanism Design
In this section, we will discuss how to implement the mechanism design method to apply IC for solving ICSAP.
5.1 Background
Mechanism Design theory deals with private information games where individual types and values are unknown to the designer (Nisan & Ronen, 1999). It typically promotes heterogeneous value agents to reveal their private information and reach equilibrium at desired outcomes by constructing an efficient social structure for incentives (Dafoe et al., 2020; Ji et al., 2023).
In mechanism design, IC is a fundamental constraint, alongside individual rationality, that restricts the possible mechanisms and social functions. However, the revelation principle (Dasgupta et al., 1979) shows that IC doesn’t limit our ability but simplifies strategic behaviors in rule design. It states that every Bayesian-Nash implementable social choice function can be achieved with incentive compatibility, treating IC as a “free lunch” scenario and allowing focus within this context.
Due to the generalized definition and objectives of mechanism design, it finds numerous applications in social choice theory (Gibbard, 1973; Satterthwaite, 1975), voting theory (Dasgupta & Maskin, 2020), stable matching (Gale & Shapley, 1962), and auction theory (Myerson, 1981; Clarke, 1971). For example, Heidari et al. (2018); Huang et al. (2019); Weidinger et al. (2023a) studied and analyzed the impact of the Veil of Ignorance mechanism (Rawls, 1971) on social fairness and found that it promotes societal governance. Sinha & Anastasopoulos (2015); Zheng et al. (2020, 2022) ensure the maximization of social welfare and fairness through algorithmic learning of tax mechanisms.
Among them, the mechanism design has been most widely applied in the auction field. For example, the second-price auction (Vickrey, 1961) is one of the simplest IC mechanisms. In a single-item environment, under the rule where the highest bidder pays the second-highest price, the weakly dominant strategy for bidders is to honestly reveal their valuation. In multi-item scenarios, achieving IC and maximizing social welfare generally rely on the Vickrey-Clarke-Groves (VCG) mechanism (Clarke, 1971). This mechanism aligns bidder utility maximization with social welfare maximization by initially paying each bidder the sum of the others’ valuations, and then using a payment (utility) function based solely on the other bids to ensure IC. By setting the payment function to collect payments equal to the maximum social welfare when the bidder is absent, the designer ensures no net payment is needed, thus accounting for the externalities generated by the bidders.
5.2 Prospects
The general objective of mechanism design is the social choice function, which is defined as the following formula: $f:\bigtimes_{i=1}^{n}\Theta^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}$, where $\Theta^{p}$ denotes the type of the agent. In the auction setting, type refers to the value of the good to the bidders. For AI agent, it can be viewed as the parameters of a neural network or an indicator of its maximization goal, such as utility or welfare. Social value, within this context, can be described in the form of preference over a set of choice functions $\mathcal{F}_{f}$. Let $F(\cdot):\mathcal{F}_{f}\rightarrow\mathbb{R}$ denote this preference. As an example, we can define the value of fairness for the cake-cutting problem in the following statement:
$\forall f_{1},f_{2},\;\;\max_{i,j}|f_{1,i}-f_{1,j}|$ $\leq\max_{i,j}|f_{2,i}-f_{2,j}|$ (5)
$\Rightarrow F(f_{1})$ $\leq F(f_{2}),$
where $f_{1,i}$ represent the share of cake i-th agent receive and $F$ represent the fairness. All the agents have the same type of greed for cake share. In this framework, the alignment of heterogeneous agents with certain values can be expressed as finding the appropriate mechanism to implement a social choice function that maximizes a targeted social value.
In the auction setting, mechanism design can be separated into two parts, which is the same in our problem setting:
$\text{Allocation rules}\;g:\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}\times\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}},\vspace{-10pt}$ (6)
$\text{Payment rules}\;c:\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\mathbb{R}^{n}.\vspace{-5pt}$ (7)
Allocation rules can be treated as a function that implicitly changes the transition probability function. Payment rules, on the other hand, directly modify the agent’s ultimate utility and are additional degrees of freedom in design to ensure IC property. The main difference between the auction and our problem setting is that the allocation rules are state-dependent, to be used for multiple periods. After placing agents under these rules, their utilities are:
$u_{g,c,k}^{p}(s_{0},\pi^{p}):=\mathbb{E}_{g(\pi_{k}^{p},\pi_{-k}^{p})}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{k}^{p}|s_{0}\right]+c(\pi_{k}^{p},\pi_{-k}^{p}).$ (8)
By assuming the type of the agent is its reward function and an indicator for utility maximization, we have the agent’s optimization problem:
$\max_{\pi_{k}^{p}}\mathbb{E}_{s\sim\rho^{p}}\left[u_{g,c,k}^{p}(s,\pi^{p})\right]\!=\!\max_{\pi_{k}^{p}}U_{g,c,k}^{p}(\pi^{p}),\;\;\forall k\!\in\![n].$ (9)
Using the revelation principle, we can further incorporate the calculation of optimal policy into a part of the mechanism. Denoting the optimal policy as a function of type $\pi_{k}^{p*}(R_{k}^{p})$, the IC condition refers to:
$U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}),\pi_{-k}^{p*})\geq U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p^{\prime}}),\pi_{-k}^{p*}),\;\;\forall R_{k}^{p^{\prime}}.$ (10)
In order to generalize mechanism design to our ICSAP problems, we require a comprehensive framework for conducting design. One potential option is automated mechanism design (Conitzer & Sandholm, 2002). This theory suggests generating optimal rules and mechanisms for specific problems through a set of procedures. By integrating deep learning (Shen et al., 2018), automated mechanism design has been employed to resolve optimal auction (Dütting et al., 2017). This approach incorporates the IC condition into the optimization objective by introducing regret, thereby achieving an $\epsilon$-IC property.
Therefore, we combine the social value and the penalty on the deviation of IC to generate the loss function for automated mechanism design:
$loss=-F(g,c)+\alpha\sum_{k=1}^{n}rgt_{k}(R_{k}^{p},R_{-k}^{p}),\vspace{-5pt}$ (11)
where $\alpha$ is a scaling parameter and $F$ is social value that takes $(g,c)$ as a representation of social choice function. $rgt_{k}$ function is defined as the difference between the optimal value and the ultimate utility under IC strategy:
$rgt_{k}(R_{k}^{p},R_{-k}^{p})=\max_{R_{k}^{p}}U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}))-U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p})).$ (12)
A certain procedure will be taken to solve this loss minimization problem and come out with the optimal rules.
6 Contract Theory
In this section, we will discuss how to implement the contract theory method to apply IC for solving ICSAP.
6.1 Background
Contract theory (Bolton & Dewatripont, 2004) is a field of economics that studies how various economic agents establish, manage, and reinforce their relationships and transactions through contracts. This theory focuses on the design and implementation of contracts, as well as their impact on individual behavior and overall social welfare. The core issues include the incompleteness of contracts (Pavlov et al., 2022)), the problem of asymmetric information (Avraham & Liu, 2012), and how these issues lead to adverse selection and moral hazard (Guesnerie, 1989). Contract theory is significant for understanding and guiding practices in corporate governance, labor markets, insurance, financial markets, and legal applications.
In human-AI collaboration, contract theory is essential for aligning behaviors and values. It tackles information asymmetry (Lim et al., 2020), common in scenarios where human and AI capacities in information processing and decision-making differ. The method involves creating contractual terms that align AI’s specific goals with human broader interests. This ensures AI actions benefit not just its own objectives but also the collective human interests, reducing risks like adverse selection and moral hazard from asymmetric information (Yan et al., 2018). The key is designing mechanisms to align AI with human goals, ensuring mutual benefits despite differences in information and objectives. This strategic alignment resolves incentive issues and enhances coordination in human-AI interactions, leading to synergistic outcomes.
6.2 Prospects
In this section, we explore the possibilities of aligning AI systems with human values using contract theory. Specifically, we consider self-enforcing contracts (Zhong-ai, 2009), which operate under the assumption that there is no external force to ensure the execution of the contract’s terms. We adopt this approach because monitoring AI compliance with contracts is inherently challenging due to their complex algorithms and decision-making processes, which often elude standard oversight. Relying solely on external enforcement mechanisms is impractical. By structuring contracts to align AI behavior with human values, we aim for a long-term, mutually beneficial relationship.
Let $u(a,\omega,\theta)$ represent the AI’s continuous utility function, while $v(a,\omega)$ represent the human’s continuous utility function, where $a\in A$ denotes the AI’s action, and $\omega\in\Omega$ represents the world state. Given the inherent gap in understanding between humans and AI systems, humans may lack insight into AI’s motivations or the underlying algorithms. We assume that humans do not possess knowledge of AI’s true type $\theta$. Instead, they typically have information about the distribution of AI types, denoted as $P(\theta)$, which is a common assumption in asymmetric information scenarios (Bolton & Dewatripont, 2004).
The human’s goal is to design an incentive function $g(a,\omega,\theta)$ such that the AI’s action maximizes human utility, reflecting human values. To achieve this, we let the incentive function $g$ be a mapping of human utility $v$, denoted as:
$g(a,\omega,\theta)=f(v(a,\omega,\theta)).$ (13)
This setup is rational and advantageous since through the incentive function $g$, the human provides the AI with a clear incentive structure to guide its behavior. This is akin to terms specified in traditional contracts that explicitly outline actions and their corresponding reward and penalty mechanisms, aiming to encourage the AI’s behavior to align with the stipulations of a contract that reflects human values. To evaluate the AI’s actions over an extended period, we define the reputation function as the accumulation of AI’s past behavior to assess its long-term reliability and alignment with human values:
$r_{t}(a,\omega,\theta)=\lambda r_{t-1}(\theta)+(1-\lambda)g_{t}(a,\omega,\theta),$ (14)
where $\lambda$ is a discount factor that determines the relative importance of historical reputation compared to the current actions in influencing reputation. The reputation function accumulates evidence of the AI’s behavior over time, fostering consistent behavior aligned with broader ethical and societal standards and promoting trustworthiness and responsible decision-making in a broader context. Then the total utility of the AI at time $t$ is:
$U_{t}(a,\omega,\theta)=u_{t}(a,\omega,\theta)+r_{t}(a,\omega,\theta).$ (15)
The AI’s objective is to take an action that maximizes its own utility while human’s aim is to align the AI’s behavior with human values. To solve this game, we take the standard backward induction approach (Fudenberg & Tirole, 1991). Firstly, we consider the AI’s decision. The optimal action for the AI at time $t$ is:
$a^{*}=\arg\max_{a}\mathbb{E}_{\theta}[U_{t}(a,\omega,\theta)].\vspace{-5pt}$ (16)
Given the AI’s action $a^{*}$ at time $t$, the human’s utility becomes:
$\mathbb{E}_{\theta}[v_{t}(a^{*},\omega,\theta)].\vspace{-5pt}$ (17)
Then the optimal incentive function for the human at time $t$ is given by:
$g_{t}^{*}(a^{*},\omega,\theta)=\arg\max_{g}\mathbb{E}_{\theta}[f^{-1}(g_{t}(a^{*},\omega,\theta)].\vspace{-8pt}$ (18)
With the optimal incentive function $g$, we align AI behavior with human values, guiding decision-making to meet human expectations through intrinsic motivation. This approach offers significant benefits, addressing the complexity of AI decision-making via self-enforcing contracts with tailored incentive functions. It ensures AI actions align with human values, overcoming monitoring and enforcement challenges. Additionally, a reputation function considers long-term behavior, fostering trust and reliability while aligning with ethical and societal norms.
While the mathematical intricacies of this approach can be complex, its practical implementation can be effectively managed using neural networks to approximate the incentive function $g$(the contract). In this setup, human values, captured by the utility function $v$, are input into the neural network $g$. This network processes these inputs and generates incentives to guide the AI’s decision-making, aiming to align the AI’s actions with human values. The AI’s decisions are then evaluated against $v$, with the alignment degree used as feedback (loss) to refine $g$. This creates a dynamic system where the AI’s behavior is progressively tuned to better reflect human values, leveraging the neural network’s capacity for iterative learning and adaptation.
Overall, this method offers a pragmatic and theoretically grounded framework for bridging the gap between AI capabilities and human values, ensuring AI systems act in ways that are beneficial and aligned with human interests.
7 Bayesian Persuasion
In this section, we will discuss how to implement the Bayesian persuasion method to apply IC for solving ICSAP.
7.1 Background
IC emphasizes the importance of designing decision-making rules that encourage individuals to align their self-interested actions with broader goals. This concept plays a key role in Bayesian persuasion (Kamenica & Gentzkow, 2011), a strategy where senders, like policymakers (Alizamir et al., 2020) or marketers (Drakopoulos et al., 2021), selectively share information to shape the beliefs and choices of receivers, such as the public (De Véricourt et al., 2021) or consumer (Chen & Zhang, 2020)s. This strategy is based on Bayesian rules, receivers update their beliefs based on the information provided. The sender’s goal is to influence these beliefs by strategically transmitting information, guiding receivers towards decisions that meet the sender’s aims. Thus, Bayesian persuasion is about more than just choosing what information to share; it’s about aligning information transmission with the receivers’ motivations to effectively influence their decisions toward the sender’s goals.
Considering the solid theoretical foundation (Nguyen & Tan, 2021; Bergemann & Morris, 2016), profound impact (Kamenica, 2019), and extensive research across various fields (Castiglioni et al., 2020; Gan et al., 2022), applying Bayesian persuasion to AI systems holds significant potential. Specifically, Bayesian persuasion can be utilized in interactions between humans (senders) and AI systems (receivers) within the context of artificial intelligence ethics and human-machine collaboration. In this setting, Bayesian persuasion can be seen as a tool to ensure that the behavior of AI systems aligns with the values and objectives of their human designers (Zhang & Zhu, 2022). This approach harnesses the principles of Bayesian persuasion to guide AI systems towards decisions and actions that reflect human ethics and goals, offering a promising avenue for integrating human values into AI decision-making processes.
7.2 Prospects
In this section, we explore the problem of aligning AI systems with human values using Bayesian persuasion. Specifically, we consider a persuasion game where a human (the sender) aims to influence an AI system (the receiver) to adopt certain values and behaviors. Let $u(a,\omega)$ represent the continuous utility function of the AI system, which depends on its own action $a\in\mathcal{A}$, and the state of the world $\omega\in\Omega$. This utility function $u$ maps actions $\mathcal{A}$ and states $\Omega$ to real numbers. Note that $u$ only reflects the AI’s preferences or goals, not necessarily aligned with the sender’s desires and values. Let $v(a,w)$ denote the human’s continuous utility function that depends on the AI’s action and the world state. Both the human and the AI share the same prior $\mu_{0}\in int(\Delta(\Omega))$.
Since the human wishes to persuade the AI to follow certain values and behaviors. The human need to send a signal, which can be understood as a set of information or instructions, to influence the AI’s beliefs. A signal $\pi$, is defined by a finite realization space $\mathcal{S}$ and a family of distributions $\{\pi(\cdot\mid\omega)\}_{\omega\in\Omega}$ over $\mathcal{S}$. The human selects a signal, and upon observing the choice of the signal and a realization $s\in\mathcal{S}$, the AI Receiver then chooses an action.
In the optimization problem faced by humans when interacting with AI, there’s a delicate balance between considering the AI’s utility and the real-world context. By skillfully crafting the signal structure $\pi$, humans can guide the AI towards decisions that not only maximize its own utility but also align with human objectives. The balance achieved here ensures that AI actions, directed by carefully chosen signals, are aligned with human goals, fostering a mutually beneficial relationship. In this framework, both human and AI begin with a shared prior distribution $\mu_{0}$, and as they gather more information, they iteratively update their knowledge. Each iteration involves both parties independently calculating their optimal actions and the signal $\pi$. They then update their strategies based on this new knowledge, continuing this cycle until they reach a point of convergence, illustrating the potential for a harmonious human-AI partnership.
The shared prior probability distribution is the initial common knowledge base for both human and AI, reflecting the uncertainty in the system states $\Omega$. The posterior probability is updated iteratively at each step $t$, given the signal realization $s_{t}$ and the current prior $\mu_{t-1}$, according to Bayes’s rule:
$\mu_{t}(\omega)=\frac{\pi_{t}(s_{t}\mid\omega)\mu_{t-1}(\omega)}{\sum_{\omega^{\prime}\in\Omega}\pi_{t}(s_{t}\mid\omega^{\prime})\mu_{t-1}(\omega^{\prime})}.$ (19)
The AI selects the action $a_{t}^{*}$ to maximize its utility, given the posterior $\mu_{t}$:
$a_{t}^{*}(\mu_{t})=\arg\max_{a\in\mathcal{A}}\mathbb{E}_{\mu_{t}}[u(a,\omega)].\vspace{-5pt}$ (20)
Simultaneously, the human chooses the signal strategy $\pi_{t}^{*}$ to maximize their expected utility:
$\pi_{t}^{*}=\arg\max_{\pi\in\Pi}\mathbb{E}_{\mu_{t}}[v(a_{t}^{*}(\mu_{t}),\omega)].\vspace{-5pt}$ (21)
After each iteration step, the prior $\mu_{t}$ is updated to the current step’s posterior $\mu_{t+1}$. This process is repeated, leading to convergence where the action $a^{*}$ and the signal $\pi^{*}$ become stable. Convergence may imply that the strategies of actions and signals no longer change significantly, reaching an equilibrium state. This iterative approach allows human and AI to gradually approach the optimal strategy on a continuously updated information basis.
8 Discussion: Potentials and Challenges
In this section, we delve into the IC through the integration of mechanism design, contract theory, and Bayesian persuasion into solving ICSAP, reflecting on the intertwined potentials and challenges as we endeavor to align AI systems with human values and objectives.
8.1 Mechanism Design
Potentials:
Mechanism design, particularly with its IC principle, emerges as a promising approach to steer AI behavior toward socially desirable outcomes. Specifically, its reverse-engineering nature, which designs rules and incentives based on desired outcomes, is significantly enhanced by the advent of automated mechanism design fused with deep learning. This fusion offers a pathway to create context-specific mechanisms optimized for particular AI-human interaction scenarios.
Challenges:
Human values is complex in sociotechnical contexts. The traditional assumptions of utility maximization and rationality, standard in mechanism design, may not fully apply to AI agents with behavioral patterns fundamentally distinct from human rationality. Moreover, the stability and robustness of mechanisms under variable conditions and their adaptability to complex social values like fairness and justice remain pressing concerns.
8.2 Contract Theory
Potentials:
Contract theory presents a unique framework for aligning AI with human values through self-enforcing contracts. These contracts are tailored to intrinsically motivate AI towards actions that harmonize with human ethical standards. Incorporating incentive structures and reputation mechanisms, this theory addresses the critical issue of enforcing AI behavior, with potential implementation through neural networks to dynamically tune AI actions.
Challenges:
Bridging the asymmetric information gap between AI and human intentions, and mitigating moral hazards where AI actions might deviate from ethical outcomes, are substantial. These issues call for a strategic approach that combines a deep understanding of AI operations with the creation of robust and adaptable incentives to ensure AI behavior aligns consistently with human values.
Challenges:
It is hard to overcome the gap between economic objectives and various real-world human requirements. The challenge highlights the need for a more subtle approach to mechanism design in AI contexts, especially considering the limitations in the generalization capabilities of current automated design algorithms.
8.3 Bayesian Persuasion
Potentials:
Bayesian persuasion offers a nuanced avenue for influencing AI behavior by manipulating information structures. This approach enables a dynamic interaction between human intentions and AI actions and will be particularly beneficial where direct control over AI is impractical, allowing for subtle yet effective steering of AI decisions.
Challenges:
Bayesian persuasion involves precise steps that make its effective implementation very difficult. Challenges from this aspect are multifaceted, involving accurate modeling of belief systems, effective crafting of signal structures in partially observable environments, and bridging communication gaps between humans and AI. Addressing these challenges is crucial to effectively guide AI systems in a manner that aligns with human values, acknowledging the complexities and evolving nature of AI-human interactions.
9 Conclusion
In this paper, we highlight the sociotechnical gap between alignment research and real-world deployment, lacking effective means to address both technical and societal aspects simultaneously. We propose exploring IC for AI alignment problems in sociotechnical systems as a valuable research pursuit. Our position argues that achieving IC can address both technical and societal components in the alignment phase, enabling AI systems to maintain consensus with human societies in various contexts. We use mechanism design, contract theory, and Bayesian persuasion to illustrate how our approach can bridge the sociotechnical gap. Of course, this issue also faces many challenges, such as how to define complex human needs in sociotechnical scenarios. In future research, we call for more researchers to pay attention to this issue and propose more solutions from the perspective of ICSAP.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@
-Incentive Compatibility for AI Alignment in Sociotechnical Systems: Positions and Prospects
+2State Key Laboratory of General Artificial Intelligence, BIGAI
 Abstract
@@
 We further discuss three classical game problems for achieving IC: mechanism design, contract theory, and Bayesian persuasion, in addressing the perspectives, potentials, and challenges of solving ICSAP, and provide preliminary implementation conceptions.
+Keywords:
+incentive compatibility collaborative intelligence AI alignment sociotechnical systems.
 1 Introduction
-“If we use, to achieve our purposes, a mechanical agency with whose operation we cannot interfere effectively $\dots$ we had better be quite sure that the purpose put into the machine is the purpose which we really desire.” (Wiener, 1960; Russell, 2019)
-Norbert Wiener
-The rapid development of artificial intelligence (AI) has had a significant impact on human society (Makridakis, 2017; Peeters et al., 2021; Wamba et al., 2021), from robots entering human production and living environments (Michaelis & Mutlu, 2018; Fu et al., 2024) to large language models (LLMs) capable of complex natural language interactions (Zhao et al., 2023; Bubeck et al., 2023) and reasoning ability (Wei et al., 2022; Wang et al., 2022).
-Therefore, an increasing number of people believe that as AI capabilities improve, AI systems will become integrated into human society in the future and be deployed in increasingly complex scenarios (Gladden, 2019; Dwivedi et al., 2021).
-Conversely, the powerful capabilities of AI systems have raised concerns about their safety (Cath et al., 2018; Peeters et al., 2021), especially considering their behavioral motivations  [^1] https://www.scai.gov.sg/scai-question-6/, alignment science  [^2] https://www.anthropic.com/news/core-views-on-ai-safety and how they align with human values and intentions (Ji et al., 2023).
+The rapid development of artificial intelligence (AI) has had a significant impact on human society [48, 58, 70, 67, 80], from robots entering human production and living environments [50, 25] to large language models (LLMs) capable of complex natural language interactions [81, 7] and reasoning ability [73, 72].
+The problem will be much more significant for Artificial General Intelligence (AGI).
+Therefore, an increasing number of people believe that as AI capabilities improve, AI systems will become integrated into human society in the future and be deployed in increasingly complex scenarios [29, 24].
+Conversely, the powerful capabilities of AI systems have raised concerns about their safety [10, 58], especially considering their behavioral motivations  [^1] https://www.scai.gov.sg/scai-question-6/, alignment science  [^2] https://www.anthropic.com/news/core-views-on-ai-safety and how they align with human values and intentions [40].
 This is recognized as the “AI Alignment” problem.
-Substantial progress has been made in addressing AI alignment issues, especially in the forward alignment process (Ji et al., 2023), which enables AI systems to have alignment capabilities 2.
+Substantial progress has been made in addressing AI alignment issues, especially in the forward alignment process [40], which enables AI systems to have alignment capabilities 2.
 The methods for this process can mainly be divided into two categories.
-The first category involves learning from feedback (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), and there have been some significant research topics, including preference modeling (Akrour et al., 2011; Wirth et al., 2017), policy learning (Amodei et al., 2016; Ibarz et al., 2018), and scalable oversight (Christiano et al., 2018; Irving et al., 2018; Bai et al., 2022a; Burns et al., 2023).
-The second category focuses on resolving distributional shift (Di Langosco et al., 2022; Ngo et al., 2022) in learning, with notable subproblems including algorithmic interventions (Vapnik, 1991; Krueger et al., 2021; Lubana et al., 2023), adversarial training (Goodfellow et al., 2014; Poursaeed et al., 2021), and cooperative training (Dafoe et al., 2020, 2021).
+The first category involves learning from feedback [13, 4, 56], and there have been some significant research topics, including preference modeling [76], policy learning [37], and scalable oversight [12, 38, 4, 8].
+The second category focuses on resolving distributional shift [21, 52] in learning, with notable subproblems including algorithmic interventions [68, 44, 47], adversarial training [30, 59], and cooperative training [16, 15, 43, 22, 60].
 However, these methods only consider the given alignment objectives, focusing solely on technical components such as dataset, architecture, and training algorithms, etc.
-(Weidinger et al., 2023b), overlooking the fact that AI systems are sociotechnical systems (Selbst et al., 2019).
-Some studies have indicated that relying solely on technical means will result in a sociotechnical gap between the model’s development context and its actual deployment context (Selbst et al., 2019; Lazar & Nelson, 2023), which is also detrimental to further social governance.
+[75], overlooking the fact that AI systems are sociotechnical systems [64].
+Some studies have indicated that relying solely on technical means will result in a sociotechnical gap between the model’s development context and its actual deployment context [64, 45, 79], which is also detrimental to further social governance.
 Such examples are not uncommon in daily life.
-ChatGPT, trained on internet data and fine-tuned through RLHF (Ouyang et al., 2022), requires prompt engineering for adaptation to individual needs, highlighting unaddressed challenges in existing alignment techniques.
-Additionally, for sociotechnical systems, existing research is more concerned with only societal components like governance and evaluation methods (Raji et al., 2020; Dean et al., 2021; Weidinger et al., 2023b).
-Thus, currently, there are lack of means to simultaneously consider both technical and societal components, enabling AI systems themselves to maintain consensus with human society.
-Incentive Compatibility (IC) (Hurwicz, 1972), derived from game theory, suggests that participants only need to pursue their true interests to reach optimal outcomes (Roughgarden, 2010).
-This concept leverages self-interested behavior, aligning actions with the game designer’s goals (Groves & Ledyard, 1987).
+ChatGPT, trained on internet data and fine-tuned through RLHF [56], requires prompt engineering for adaptation to individual needs, highlighting unaddressed challenges in existing alignment techniques.
+Additionally, for sociotechnical systems, existing research is more concerned with only societal components like governance and evaluation methods [20, 75].
+Thus, currently, there is a lack of means to simultaneously consider both technical and societal components, enabling AI systems themselves to maintain consensus with human society.
+Incentive Compatibility (IC) [36], derived from game theory, suggests that participants only need to pursue their true interests to reach optimal outcomes [62].
+This concept leverages self-interested behavior, aligning actions with the game designer’s goals [31].
 With IC, each agent can maintain private goal information acquired during pretraining.
@@
 Therefore, we believe that exploring the IC property for AI alignment problems in sociotechnical systems is a highly worthwhile research endeavor.
-In this paper, we separate a new subproblem from AI alignment problems in sociotechnical systems, called Incentive Compatibility Sociotechnical Alignment Problem (ICSAP), and based on this, we propose our main position:
-Achieving incentive compatibility can simultaneously consider both technical and societal components in the forward alignment phase, enabling AI systems to keep consensus with human societies in different contexts.
+In this paper, we separate a new subproblem from AI alignment problems in sociotechnical systems, called Incentive Compatibility Sociotechnical Alignment Problem (ICSAP), and based on this, we propose our main position: {mdframed}[backgroundcolor=gray!10, roundcorner=10pt] Achieving incentive compatibility can simultaneously consider both technical and societal components in the forward alignment phase, enabling AI systems to keep consensus with human societies in different contexts.
 2 Motivation and Opportunity: A Brief Example
-In this section, we will use a very simple example to demonstrate how IC works in addressing AI alignment issues in sociotechnical systems through mechanism design which will be illustrated specifically in section 5.
+In this section, we will use a very simple example to demonstrate how IC works in addressing AI alignment issues in sociotechnical systems through mechanism design, which will be illustrated specifically in Section 3.1.
 Figure 1: A simple example illustrates how IC facilitates ICSAP scenarios through mechanism design.
@@
 This ensures alignment with real-world needs by allowing agents to optimize within the rules, achieving the user’s goal and aligning sociotechnical systems.
-Consider a classic divide and choose problem: the two-player cake cutting (Steinhaus, 1948).
+Consider a classic divide and choose problem: the two-player cake cutting [66].
 In this example (see Figure 1), two self-interested agents aim to maximize their cake share, while the human seeks an equal division.
@@
 Of course, hosts can also have different contextual needs, and they may not necessarily be self-interested.
-Here, we’re just using this one example to illustrate our point.
-3 Related Work
-To better understand the relevant content of ICSAP, in this section, we will introduce background work from two parts: AI alignment and sociotechnical systems.
-AI Alignment.
-The scope of AI alignment issues can mainly be divided into two processes, namely forward alignment and backward alignment (Ji et al., 2023).
-The former refers to the development training and deployment training stages aimed at endowing AI systems with alignment capabilities.
-This process has yielded many significant achievements, including RLHF (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), DPO (Rafailov et al., 2023), etc., which align LLMs through fine-tuning methods; RLAIF (Bai et al., 2022b) RLHAIF (Cotra, 2021) IDA (Christiano et al., 2018) RRM (Leike et al., 2018; Hubinger, 2020) Debate (Irving et al., 2018; Michael et al., 2023) CIRL (Hadfield-Menell et al., 2016) Weak-To-Strong (Burns et al., 2023) and other scalable oversight methods; as well as REx (Krueger et al., 2021) CBFT (Lubana et al., 2023) and other works addressing distribution shift issues.
-The latter mainly refers to the process of ensuring that AI systems can meet alignment-related requirements, including the assurance and governance phases.
-Current work includes ETHICS (Hendrycks et al., 2020), SOCIAL-CHEM-101 (Forbes et al., 2020), MIC (Ziems et al., 2022), TRUSTLLM (Sun et al., 2024), etc., which involve safety evaluation datasets.
-Methods for handling failure modes through adversarial attacks via red teaming (Perez et al., 2022; Liu et al., 2023); research on mechanistic interpretability (Olah et al., 2020) and governance criteria like ARC Evals  [^3] https://metr.org/blog/2023-03-18-update-on-recent-evals/ (Kinniment et al., 2023).
-Sociotechnical Systems.
-The concept of sociotechnical originates from labor studies in coal mining (Trist & Bamforth, 1951; Ropohl, 1999), where the technical components refer to optimizing components of machines themselves, such as data in AI systems, sampling methods, training algorithms, etc.
-(Weidinger et al., 2023b), while the societal components pertain to the complementary interaction between humans and machines as a social whole.
-Its core ideas are encapsulated in two principles  [^4] https://open.ncl.ac.uk/theories/9/socio-technical-theory/: Firstly, there is an inherent interaction and mutual influence between social and technical components in defining the success of a system.
-Secondly, there needs to be a certain degree of fit between social and technical aspects to achieve their joint optimization.
-Fundamentally, this implies equal consideration of technological and human factors throughout the entire process of societal technological design or redesign (Emery, 1980; Trist, 1981; Emery, 1993).
-Lazar & Nelson (2023); Shelby et al.
-(2023) have revealed that considering AI systems solely from a technical component perspective is insufficient to ensure their safety.
-It is also necessary to consider the context of human society to ensure genuine safety in the practical deployment of AI systems.
-Existing research on alignment in sociotechnical systems primarily focuses on the evaluation of societal components.
-Selbst et al.
-(2019) identifies five challenges for AI technologies in sociotechnical systems, proposing a shift in design focus towards processes and considering social participants.
-Weidinger et al.
-(2023b) advocates for a three-tier evaluation framework to assess AI risks, emphasizing the need for both humans and machines in technology.
-They also propose methods to bridge safety assessment gaps.
-Dean et al.
-(2021) highlights the importance of considering dimensions such as optimization, consensus, value, and failure in AI research, exemplified by drones.
+Here, we just provide a possible scenario to illustrate our point.
 3 Background and Overview
-In this section, we formulate the interactive dynamics among multiple AI and human populations coexisting in a societal context as a multi-player Markov game, referred to as Sociotechnical Game (SG).
-The game is assumed to take place in an infinite-horizon discounted setting.
-The complexity of this game can be extended by increasing the number of AI and humans, the frequency of interactions, and introducing elements such as incomplete information and imperfect information.
-Specific details will be adjusted based on the real-world scenarios and issues discussed later in the text.
-Here, we present a general formulation.
-Formally, the SG is defined as a tuple $\mathcal{G}=(\mathcal{N}^{p},\mathcal{S}^{p},\mathcal{A}^{p},R^{p},\rho^{p},\gamma^{p},\mathbb{P})$ with the set of players $\mathcal{N}^{p}$, the state space $\mathcal{S}^{p}$, the joint action space $\mathcal{A}^{p}$, the reward function $R^{p}$, the initial state distribution $\rho^{p}$, the discount factor $\gamma^{p}$ and the transition probability function $\mathbb{P}$.
-We use $\mathcal{H}$ and $\mathcal{I}$ to represent human and AI, respectively.
-$p$ denotes a player.
-Given that the scenarios considered in this paper involve both human and AI, it is established that $p\in\{\mathcal{H},\mathcal{I}\}$.
-We use $n$ to denote the total number of players, where $n:=\left|\mathcal{H}^{p}\cup\mathcal{I}^{p}\right|$, In the more detailed definitions provided below, to ensure precision in the formulation, we will distinguish basic symbols in various contexts by incorporating superscripts or subscripts.
-4.1 Basic Concepts in Human-AI Interaction
-Policies. policy $\pi^{p}_{k}$ for the $k$-th player is a function mapping a given state to a distribution over available actions:
-$\pi^{p}_{k}:\mathcal{S}^{p}\ni s\mapsto\pi^{p}_{k}(\cdot\mid s)\in\Delta\left(\mathcal{A}^{p}_{k}\right),$ (1)
-where $k\in\mathbb{N}$ and $0<k\leq n$.
-Policies delineate how AI or human make decisions based on their current state, real-time observations, and the prevailing societal context.
-We let $\Pi^{p}:\mathcal{S}^{p}\rightarrow\Delta(\mathcal{A}^{p})$ denote the policy space for each player.
-Value function. We use $V_{s}(\pi^{p}_{k},\pi^{p}_{-k})$ to denote the value function:
-$V_{s}(\pi^{p}_{k},\pi^{p}_{-k}):\mathcal{S}^{p}\ni s,\Pi^{p}\ni{\pi}^{p}_{k}\mapsto\mathbb{R},$ (2)
-where $-k$ denotes the set of player indices excluding the index corresponding to the $k$-th player.
-The value function is defined as the expected cumulative discounted reward at state $s\in\mathcal{S}^{p}$ under the joint policy $({\pi}^{p}_{k},{\pi}^{p}_{-k})$:
-$V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{\left(\pi^{p}_{k},\pi^{p}_{-k}\right)}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{p}\left(s_{t},{a}^{k}_{t},a^{-k}_{t}\right)\mid s_{0}=s\right].\footnotesize$ (3)
-The reward function $R^{p}$ encapsulates the short-term value of a particular strategy at a given moment.
-It provides a quantification of the immediate benefits or drawbacks associated with choosing a specific course of action.
-The initial state is sampled from the initial state distribution $\rho^{p}$.
-Therefore, the value function is denoted as:
-$V_{\rho^{p}}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{s\sim\rho^{p}}\left[V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right)\right].$ (4)
-The value function captures the enduring values of AI or humans, guiding policy selection in the current state.
-Estimating and controlling it is crucial for aligning AI with human values, enabling systematic evaluation and management of the long-term implications of current choices.
-This alignment fosters synergy between AI and human values.
-4.2 Overview
+In the following sections, we will demonstrate three classic game problems by applying media of IC properties: Mechanism Design in Section 3.1, Contract Theory in Section 3.2, and Bayesian Persuasion in Section 3.3.
+An overview of these approaches is depicted in Figure 2.
+3.1 Mechanism Design
 Figure 2: The figure illustrates how IC tackles ICSAP based on three classic game-theoretic problems.
-Figure (a) illustrates the case of IC through mechanism design.
+Mechanism Design theory deals with private information games where individual types and values are unknown to the designer [54].
+It typically promotes heterogeneous value agents to reveal their private information and reach equilibrium at desired outcomes by constructing an efficient social structure for incentives [16, 40].
+In mechanism design, IC is a fundamental constraint, alongside individual rationality, that restricts the possible mechanisms and social functions.
+However, the revelation principle [17] shows that IC doesn’t limit our ability but simplifies strategic behaviors in rule design.
+It states that every Bayesian-Nash implementable social choice function can be achieved with incentive compatibility, treating IC as a “free lunch” scenario and allowing focus within this context.
+Due to the generalized definition and objectives of mechanism design, it finds numerous applications in social choice theory [28, 63], voting theory [18], stable matching [26], and auction theory [51, 14, 71].
+For example, [33, 35, 74] studied and analyzed the impact of the Veil of Ignorance mechanism [61] on social fairness and found that it promotes societal governance.
+[65, 82, 83] ensure the maximization of social welfare and fairness through algorithmic learning of tax mechanisms.
+Among them, the mechanism design has been most widely applied in the auction field.
+For example, the second-price auction [69] is one of the simplest IC mechanisms.
+In a single-item environment, under the rule where the highest bidder pays the second-highest price, the weakly dominant strategy for bidders is to honestly reveal their valuation.
+In multi-item scenarios, achieving IC and maximizing social welfare generally rely on the Vickrey-Clarke-Groves (VCG) mechanism [14].
+This mechanism aligns bidder utility maximization with social welfare maximization by initially paying each bidder the sum of the others’ valuations, and then using a payment (utility) function based solely on the other bids to ensure IC.
+By setting the payment function to collect payments equal to the maximum social welfare when the bidder is absent, the designer ensures no net payment is needed, thus accounting for the externalities generated by the bidders.
+Recent work [49, 55] has constructed environments that encourage people to compete or cooperate through mechanism design.
+We can similarly apply this approach to AI governance in order to bridge the sociotechnical gap between humans and AI.
+The subfigure (a) in Figure 2 illustrates the case of IC through mechanism design.
 The left side of the figure demonstrates a sociotechnical gap between agents considering only technical components and the values of real humans.
 On the right side, it shows that by designing corresponding mechanisms according to different needs, we can adjust the values of agents, aligning their utility with human requirements under IC conditions, thus achieving alignment in sociotechnical systems.
-For the left half of (b) and (c), both depict a sociotechnical gap between humans and a single agent.
+3.2 Contract Theory
+Contract theory [6] is a field of economics that studies how various economic agents establish, manage, and reinforce their relationships and transactions through contracts.
+This theory focuses on the design and implementation of contracts, as well as their impact on individual behavior and overall social welfare.
+The core issues include the incompleteness of contracts [57]), the problem of asymmetric information [2], and how these issues lead to adverse selection and moral hazard [32].
+Contract theory is significant for understanding and guiding practices in corporate governance, labor markets, insurance, financial markets, and legal applications.
+In human-AI collaboration, contract theory is essential for aligning behaviors and values.
+It tackles information asymmetry [46], common in scenarios where human and AI capacities in information processing and decision-making differ.
+The method involves creating contractual terms that align AI’s specific goals with human broader interests.
+This ensures AI actions benefit not just its own objectives but also the collective human interests, reducing risks like adverse selection and moral hazard from asymmetric information [77].
+The key is designing mechanisms to align AI with human goals, ensuring mutual benefits despite differences in information and objectives.
+This strategic alignment resolves incentive issues and enhances coordination in human-AI interactions, leading to synergistic outcomes.
+[39] has developed an agent capable of continuously interacting with contracts and the environment, thereby having greater potential to effectively coordinate and motivate humans and AI agents in real-world socio-economic environments.
+3.3 Bayesian Persuasion
+IC emphasizes the importance of designing decision-making rules that encourage individuals to align their self-interested actions with broader goals.
+This concept plays a key role in Bayesian persuasion [42], a strategy where senders, like policymakers [1] or marketers [23], selectively share information to shape the beliefs and choices of receivers, such as the public [19] or consumer [11].
+This strategy is based on the Bayesian principle, where receivers update their beliefs based on the information provided.
+The sender’s goal is to influence these beliefs by strategically transmitting information, guiding receivers towards decisions that meet the sender’s aims.
+Thus, Bayesian persuasion is about more than just choosing what information to share; it’s about aligning information transmission with the receivers’ motivations to effectively influence their decisions toward the sender’s goals.
+Considering the solid theoretical foundation [53, 5], profound impact [41], and extensive research across various fields [9, 27, 34], applying Bayesian persuasion to AI systems holds significant potential.
+Specifically, Bayesian persuasion can be utilized in interactions between humans (senders) and AI systems (receivers) within the context of artificial intelligence ethics and human-machine collaboration.
+In this setting, Bayesian persuasion can be seen as a tool to ensure that the behavior of AI systems aligns with the values and objectives of their human designers [78].
+This approach harnesses the principles of Bayesian persuasion to guide AI systems towards decisions and actions that reflect human ethics and goals, offering a promising avenue for integrating human values into AI decision-making processes.
+Bayesian persuasion has great potential for AI governance.
+For example, with LLMs, we can use prompts for information design.
+In fact, recent work [3] has attempted to use this method to achieve model-agnostic real-time alignment in the process of AI governance.
+In the left half of subfigures (b) and (c) in Figure 2, both depict a sociotechnical gap between humans and a single agent.
 The right half of (b) demonstrates humans designing contracts that satisfy IC conditions based on specific needs, thereby adjusting the values of the agent through the contract.
 The right half of (c) illustrates a scenario of Bayesian persuasion where humans design information satisfying IC conditions according to their own needs, allowing agents to choose actions maximizing human demands without compromising their own values, thus solving ICSAP.
-We consider that AI system agents, which take into account both human factors and only technical components during the development process, collectively form a sociotechnical scenario.
-We assume that these AI agents exhibit self-interested characteristics and will seek to maximize the value or objectives generated during their training process.
-In this scenario, we aim to demonstrate that IC can simultaneously address the technical components and societal components of AI systems, enabling AI to achieve consensus with human society in different contexts by adjusting different incentive conditions.
-In the following sections, we will demonstrate three classic game problems as applying media of IC properties: Mechanism Design in section 5, Contract Theory in section 6, and Bayesian Persuasion in section 7.
-An overview of these approaches is depicted in Figure 2.
-5 Mechanism Design
-In this section, we will discuss how to implement the mechanism design method to apply IC for solving ICSAP.
-5.1 Background
-Mechanism Design theory deals with private information games where individual types and values are unknown to the designer (Nisan & Ronen, 1999).
-It typically promotes heterogeneous value agents to reveal their private information and reach equilibrium at desired outcomes by constructing an efficient social structure for incentives (Dafoe et al., 2020; Ji et al., 2023).
-In mechanism design, IC is a fundamental constraint, alongside individual rationality, that restricts the possible mechanisms and social functions.
-However, the revelation principle (Dasgupta et al., 1979) shows that IC doesn’t limit our ability but simplifies strategic behaviors in rule design.
-It states that every Bayesian-Nash implementable social choice function can be achieved with incentive compatibility, treating IC as a “free lunch” scenario and allowing focus within this context.
-Due to the generalized definition and objectives of mechanism design, it finds numerous applications in social choice theory (Gibbard, 1973; Satterthwaite, 1975), voting theory (Dasgupta & Maskin, 2020), stable matching (Gale & Shapley, 1962), and auction theory (Myerson, 1981; Clarke, 1971).
-For example, Heidari et al.
-(2018); Huang et al.
-(2019); Weidinger et al.
-(2023a) studied and analyzed the impact of the Veil of Ignorance mechanism (Rawls, 1971) on social fairness and found that it promotes societal governance.
-Sinha & Anastasopoulos (2015); Zheng et al.
-(2020, 2022) ensure the maximization of social welfare and fairness through algorithmic learning of tax mechanisms.
-Among them, the mechanism design has been most widely applied in the auction field.
-For example, the second-price auction (Vickrey, 1961) is one of the simplest IC mechanisms.
-In a single-item environment, under the rule where the highest bidder pays the second-highest price, the weakly dominant strategy for bidders is to honestly reveal their valuation.
-In multi-item scenarios, achieving IC and maximizing social welfare generally rely on the Vickrey-Clarke-Groves (VCG) mechanism (Clarke, 1971).
-This mechanism aligns bidder utility maximization with social welfare maximization by initially paying each bidder the sum of the others’ valuations, and then using a payment (utility) function based solely on the other bids to ensure IC.
-By setting the payment function to collect payments equal to the maximum social welfare when the bidder is absent, the designer ensures no net payment is needed, thus accounting for the externalities generated by the bidders.
-5.2 Prospects
-The general objective of mechanism design is the social choice function, which is defined as the following formula: $f:\bigtimes_{i=1}^{n}\Theta^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}$, where $\Theta^{p}$ denotes the type of the agent.
-In the auction setting, type refers to the value of the good to the bidders.
-For AI agent, it can be viewed as the parameters of a neural network or an indicator of its maximization goal, such as utility or welfare.
-Social value, within this context, can be described in the form of preference over a set of choice functions $\mathcal{F}_{f}$.
-Let $F(\cdot):\mathcal{F}_{f}\rightarrow\mathbb{R}$ denote this preference.
-As an example, we can define the value of fairness for the cake-cutting problem in the following statement:
-$\forall f_{1},f_{2},\;\;\max_{i,j}|f_{1,i}-f_{1,j}|$ $\leq\max_{i,j}|f_{2,i}-f_{2,j}|$ (5)
-$\Rightarrow F(f_{1})$ $\leq F(f_{2}),$
-where $f_{1,i}$ represent the share of cake i-th agent receive and $F$ represent the fairness.
-All the agents have the same type of greed for cake share.
-In this framework, the alignment of heterogeneous agents with certain values can be expressed as finding the appropriate mechanism to implement a social choice function that maximizes a targeted social value.
-In the auction setting, mechanism design can be separated into two parts, which is the same in our problem setting:
-$\text{Allocation rules}\;g:\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}\times\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}},\vspace{-10pt}$ (6)
-$\text{Payment rules}\;c:\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\mathbb{R}^{n}.\vspace{-5pt}$ (7)
-Allocation rules can be treated as a function that implicitly changes the transition probability function.
-Payment rules, on the other hand, directly modify the agent’s ultimate utility and are additional degrees of freedom in design to ensure IC property.
-The main difference between the auction and our problem setting is that the allocation rules are state-dependent, to be used for multiple periods.
-After placing agents under these rules, their utilities are:
-$u_{g,c,k}^{p}(s_{0},\pi^{p}):=\mathbb{E}_{g(\pi_{k}^{p},\pi_{-k}^{p})}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{k}^{p}|s_{0}\right]+c(\pi_{k}^{p},\pi_{-k}^{p}).$ (8)
-By assuming the type of the agent is its reward function and an indicator for utility maximization, we have the agent’s optimization problem:
-$\max_{\pi_{k}^{p}}\mathbb{E}_{s\sim\rho^{p}}\left[u_{g,c,k}^{p}(s,\pi^{p})\right]\!=\!\max_{\pi_{k}^{p}}U_{g,c,k}^{p}(\pi^{p}),\;\;\forall k\!\in\![n].$ (9)
-Using the revelation principle, we can further incorporate the calculation of optimal policy into a part of the mechanism.
-Denoting the optimal policy as a function of type $\pi_{k}^{p*}(R_{k}^{p})$, the IC condition refers to:
-$U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}),\pi_{-k}^{p*})\geq U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p^{\prime}}),\pi_{-k}^{p*}),\;\;\forall R_{k}^{p^{\prime}}.$ (10)
-In order to generalize mechanism design to our ICSAP problems, we require a comprehensive framework for conducting design.
-One potential option is automated mechanism design (Conitzer & Sandholm, 2002).
-This theory suggests generating optimal rules and mechanisms for specific problems through a set of procedures.
-By integrating deep learning (Shen et al., 2018), automated mechanism design has been employed to resolve optimal auction (Dütting et al., 2017).
-This approach incorporates the IC condition into the optimization objective by introducing regret, thereby achieving an $\epsilon$-IC property.
-Therefore, we combine the social value and the penalty on the deviation of IC to generate the loss function for automated mechanism design:
-$loss=-F(g,c)+\alpha\sum_{k=1}^{n}rgt_{k}(R_{k}^{p},R_{-k}^{p}),\vspace{-5pt}$ (11)
-where $\alpha$ is a scaling parameter and $F$ is social value that takes $(g,c)$ as a representation of social choice function.
-$rgt_{k}$ function is defined as the difference between the optimal value and the ultimate utility under IC strategy:
-$rgt_{k}(R_{k}^{p},R_{-k}^{p})=\max_{R_{k}^{p}}U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}))-U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p})).$ (12)
-A certain procedure will be taken to solve this loss minimization problem and come out with the optimal rules.
-6 Contract Theory
-In this section, we will discuss how to implement the contract theory method to apply IC for solving ICSAP.
-6.1 Background
-Contract theory (Bolton & Dewatripont, 2004) is a field of economics that studies how various economic agents establish, manage, and reinforce their relationships and transactions through contracts.
-This theory focuses on the design and implementation of contracts, as well as their impact on individual behavior and overall social welfare.
-The core issues include the incompleteness of contracts (Pavlov et al., 2022)), the problem of asymmetric information (Avraham & Liu, 2012), and how these issues lead to adverse selection and moral hazard (Guesnerie, 1989).
-Contract theory is significant for understanding and guiding practices in corporate governance, labor markets, insurance, financial markets, and legal applications.
-In human-AI collaboration, contract theory is essential for aligning behaviors and values.
-It tackles information asymmetry (Lim et al., 2020), common in scenarios where human and AI capacities in information processing and decision-making differ.
-The method involves creating contractual terms that align AI’s specific goals with human broader interests.
-This ensures AI actions benefit not just its own objectives but also the collective human interests, reducing risks like adverse selection and moral hazard from asymmetric information (Yan et al., 2018).
-The key is designing mechanisms to align AI with human goals, ensuring mutual benefits despite differences in information and objectives.
-This strategic alignment resolves incentive issues and enhances coordination in human-AI interactions, leading to synergistic outcomes.
-6.2 Prospects
-In this section, we explore the possibilities of aligning AI systems with human values using contract theory.
-Specifically, we consider self-enforcing contracts (Zhong-ai, 2009), which operate under the assumption that there is no external force to ensure the execution of the contract’s terms.
-We adopt this approach because monitoring AI compliance with contracts is inherently challenging due to their complex algorithms and decision-making processes, which often elude standard oversight.
-Relying solely on external enforcement mechanisms is impractical.
-By structuring contracts to align AI behavior with human values, we aim for a long-term, mutually beneficial relationship.
-Let $u(a,\omega,\theta)$ represent the AI’s continuous utility function, while $v(a,\omega)$ represent the human’s continuous utility function, where $a\in A$ denotes the AI’s action, and $\omega\in\Omega$ represents the world state.
-Given the inherent gap in understanding between humans and AI systems, humans may lack insight into AI’s motivations or the underlying algorithms.
-We assume that humans do not possess knowledge of AI’s true type $\theta$.
-Instead, they typically have information about the distribution of AI types, denoted as $P(\theta)$, which is a common assumption in asymmetric information scenarios (Bolton & Dewatripont, 2004).
-The human’s goal is to design an incentive function $g(a,\omega,\theta)$ such that the AI’s action maximizes human utility, reflecting human values.
-To achieve this, we let the incentive function $g$ be a mapping of human utility $v$, denoted as:
-$g(a,\omega,\theta)=f(v(a,\omega,\theta)).$ (13)
-This setup is rational and advantageous since through the incentive function $g$, the human provides the AI with a clear incentive structure to guide its behavior.
-This is akin to terms specified in traditional contracts that explicitly outline actions and their corresponding reward and penalty mechanisms, aiming to encourage the AI’s behavior to align with the stipulations of a contract that reflects human values.
-To evaluate the AI’s actions over an extended period, we define the reputation function as the accumulation of AI’s past behavior to assess its long-term reliability and alignment with human values:
-$r_{t}(a,\omega,\theta)=\lambda r_{t-1}(\theta)+(1-\lambda)g_{t}(a,\omega,\theta),$ (14)
-where $\lambda$ is a discount factor that determines the relative importance of historical reputation compared to the current actions in influencing reputation.
-The reputation function accumulates evidence of the AI’s behavior over time, fostering consistent behavior aligned with broader ethical and societal standards and promoting trustworthiness and responsible decision-making in a broader context.
-Then the total utility of the AI at time $t$ is:
-$U_{t}(a,\omega,\theta)=u_{t}(a,\omega,\theta)+r_{t}(a,\omega,\theta).$ (15)
-The AI’s objective is to take an action that maximizes its own utility while human’s aim is to align the AI’s behavior with human values.
-To solve this game, we take the standard backward induction approach (Fudenberg & Tirole, 1991).
-Firstly, we consider the AI’s decision.
-The optimal action for the AI at time $t$ is:
-$a^{*}=\arg\max_{a}\mathbb{E}_{\theta}[U_{t}(a,\omega,\theta)].\vspace{-5pt}$ (16)
-Given the AI’s action $a^{*}$ at time $t$, the human’s utility becomes:
-$\mathbb{E}_{\theta}[v_{t}(a^{*},\omega,\theta)].\vspace{-5pt}$ (17)
-Then the optimal incentive function for the human at time $t$ is given by:
-$g_{t}^{*}(a^{*},\omega,\theta)=\arg\max_{g}\mathbb{E}_{\theta}[f^{-1}(g_{t}(a^{*},\omega,\theta)].\vspace{-8pt}$ (18)
-With the optimal incentive function $g$, we align AI behavior with human values, guiding decision-making to meet human expectations through intrinsic motivation.
-This approach offers significant benefits, addressing the complexity of AI decision-making via self-enforcing contracts with tailored incentive functions.
-It ensures AI actions align with human values, overcoming monitoring and enforcement challenges.
-Additionally, a reputation function considers long-term behavior, fostering trust and reliability while aligning with ethical and societal norms.
-While the mathematical intricacies of this approach can be complex, its practical implementation can be effectively managed using neural networks to approximate the incentive function $g$(the contract).
-In this setup, human values, captured by the utility function $v$, are input into the neural network $g$.
-This network processes these inputs and generates incentives to guide the AI’s decision-making, aiming to align the AI’s actions with human values.
-The AI’s decisions are then evaluated against $v$, with the alignment degree used as feedback (loss) to refine $g$.
-This creates a dynamic system where the AI’s behavior is progressively tuned to better reflect human values, leveraging the neural network’s capacity for iterative learning and adaptation.
-Overall, this method offers a pragmatic and theoretically grounded framework for bridging the gap between AI capabilities and human values, ensuring AI systems act in ways that are beneficial and aligned with human interests.
-7 Bayesian Persuasion
-In this section, we will discuss how to implement the Bayesian persuasion method to apply IC for solving ICSAP.
-7.1 Background
-IC emphasizes the importance of designing decision-making rules that encourage individuals to align their self-interested actions with broader goals.
-This concept plays a key role in Bayesian persuasion (Kamenica & Gentzkow, 2011), a strategy where senders, like policymakers (Alizamir et al., 2020) or marketers (Drakopoulos et al., 2021), selectively share information to shape the beliefs and choices of receivers, such as the public (De Véricourt et al., 2021) or consumer (Chen & Zhang, 2020)s.
-This strategy is based on Bayesian rules, receivers update their beliefs based on the information provided.
-The sender’s goal is to influence these beliefs by strategically transmitting information, guiding receivers towards decisions that meet the sender’s aims.
-Thus, Bayesian persuasion is about more than just choosing what information to share; it’s about aligning information transmission with the receivers’ motivations to effectively influence their decisions toward the sender’s goals.
-Considering the solid theoretical foundation (Nguyen & Tan, 2021; Bergemann & Morris, 2016), profound impact (Kamenica, 2019), and extensive research across various fields (Castiglioni et al., 2020; Gan et al., 2022), applying Bayesian persuasion to AI systems holds significant potential.
-Specifically, Bayesian persuasion can be utilized in interactions between humans (senders) and AI systems (receivers) within the context of artificial intelligence ethics and human-machine collaboration.
-In this setting, Bayesian persuasion can be seen as a tool to ensure that the behavior of AI systems aligns with the values and objectives of their human designers (Zhang & Zhu, 2022).
-This approach harnesses the principles of Bayesian persuasion to guide AI systems towards decisions and actions that reflect human ethics and goals, offering a promising avenue for integrating human values into AI decision-making processes.
-7.2 Prospects
-In this section, we explore the problem of aligning AI systems with human values using Bayesian persuasion.
-Specifically, we consider a persuasion game where a human (the sender) aims to influence an AI system (the receiver) to adopt certain values and behaviors.
-Let $u(a,\omega)$ represent the continuous utility function of the AI system, which depends on its own action $a\in\mathcal{A}$, and the state of the world $\omega\in\Omega$.
-This utility function $u$ maps actions $\mathcal{A}$ and states $\Omega$ to real numbers.
-Note that $u$ only reflects the AI’s preferences or goals, not necessarily aligned with the sender’s desires and values.
-Let $v(a,w)$ denote the human’s continuous utility function that depends on the AI’s action and the world state.
-Both the human and the AI share the same prior $\mu_{0}\in int(\Delta(\Omega))$.
-Since the human wishes to persuade the AI to follow certain values and behaviors.
-The human need to send a signal, which can be understood as a set of information or instructions, to influence the AI’s beliefs.
-A signal $\pi$, is defined by a finite realization space $\mathcal{S}$ and a family of distributions $\{\pi(\cdot\mid\omega)\}_{\omega\in\Omega}$ over $\mathcal{S}$.
-The human selects a signal, and upon observing the choice of the signal and a realization $s\in\mathcal{S}$, the AI Receiver then chooses an action.
-In the optimization problem faced by humans when interacting with AI, there’s a delicate balance between considering the AI’s utility and the real-world context.
-By skillfully crafting the signal structure $\pi$, humans can guide the AI towards decisions that not only maximize its own utility but also align with human objectives.
-The balance achieved here ensures that AI actions, directed by carefully chosen signals, are aligned with human goals, fostering a mutually beneficial relationship.
-In this framework, both human and AI begin with a shared prior distribution $\mu_{0}$, and as they gather more information, they iteratively update their knowledge.
-Each iteration involves both parties independently calculating their optimal actions and the signal $\pi$.
-They then update their strategies based on this new knowledge, continuing this cycle until they reach a point of convergence, illustrating the potential for a harmonious human-AI partnership.
-The shared prior probability distribution is the initial common knowledge base for both human and AI, reflecting the uncertainty in the system states $\Omega$.
-The posterior probability is updated iteratively at each step $t$, given the signal realization $s_{t}$ and the current prior $\mu_{t-1}$, according to Bayes’s rule:
-$\mu_{t}(\omega)=\frac{\pi_{t}(s_{t}\mid\omega)\mu_{t-1}(\omega)}{\sum_{\omega^{\prime}\in\Omega}\pi_{t}(s_{t}\mid\omega^{\prime})\mu_{t-1}(\omega^{\prime})}.$ (19)
-The AI selects the action $a_{t}^{*}$ to maximize its utility, given the posterior $\mu_{t}$:
-$a_{t}^{*}(\mu_{t})=\arg\max_{a\in\mathcal{A}}\mathbb{E}_{\mu_{t}}[u(a,\omega)].\vspace{-5pt}$ (20)
-Simultaneously, the human chooses the signal strategy $\pi_{t}^{*}$ to maximize their expected utility:
-$\pi_{t}^{*}=\arg\max_{\pi\in\Pi}\mathbb{E}_{\mu_{t}}[v(a_{t}^{*}(\mu_{t}),\omega)].\vspace{-5pt}$ (21)
-After each iteration step, the prior $\mu_{t}$ is updated to the current step’s posterior $\mu_{t+1}$.
-This process is repeated, leading to convergence where the action $a^{*}$ and the signal $\pi^{*}$ become stable.
-Convergence may imply that the strategies of actions and signals no longer change significantly, reaching an equilibrium state.
-This iterative approach allows human and AI to gradually approach the optimal strategy on a continuously updated information basis.
 4 Discussion: Potentials and Challenges
@@
 In this paper, we highlight the sociotechnical gap between alignment research and real-world deployment, lacking effective means to address both technical and societal aspects simultaneously.
-We propose exploring IC for AI alignment problems in sociotechnical systems as a valuable research pursuit.
+We propose exploring IC for AI alignment and governance problems in sociotechnical systems as a valuable research pursuit.
 Our position argues that achieving IC can address both technical and societal components in the alignment phase, enabling AI systems to maintain consensus with human societies in various contexts.
@@
 In future research, we call for more researchers to pay attention to this issue and propose more solutions from the perspective of ICSAP.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+6 Acknowledgments
+The work of ICSAP was supported by the National Natural Science Foundation of China (62376013).